. According to one survey in India, 75% of Instagram users love REELS. Suppose that 25
Instagram users (randomly selected) have been approached in the university located in vile
parle. They have been asked about their status of like/ dislike the Instagram- REELS.
a) What is the probability that Exactly 15 of them would agree with the claim (or said they
love Insta-REELS)?
b) What is the probability that Exactly 20 of them would agree with the claim (or said they
love Insta-REELS)?

Answer by Boreal(15235)

a. this is 25C15*0.75^15*0.25^10=0.0417
b. this is 25C20*0.75^20*0.25^5=0.1645, higher because the expected value is 25*0.75=18.75